OpenStudy (anonymous):

is by answer right? Solve. 64^x = 16^(x–1) answer? x = –2

OpenStudy (anonymous):

64x=16x-16 to 48x= -16 x = -3

OpenStudy (lgbasallote):

\[64^x = 16^{x-1}\] \[4^{3x} = 4^{2(x-1)}\] \[3x = 2(x-1)\] \[3x = 2x - 2\] \[3x - 2x = -2\] \]x = -2\]

OpenStudy (anonymous):

would that be like plugging in-1/3??

OpenStudy (anonymous):

you got correct answer